# Landau.bib

### v1

@@ 3.2 Chain rule
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\tilde{\nabla}\cdot\mathcal{U}_{t}=0.$
@@
Let $\gamma\in(-3,1]$. Let $(\mu_{t},\mathcal{U}_{t})\in\mathcal{TGRE}_{T}$ such that $|\mathcal{H}(\mu_{0})|<+\infty$ and $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. We assume
@@
    for some $p>\frac{3}{3-|\gamma|}$.
@@
    $\text{and}\quad f\in L^{\infty}([0,T];{L^{p}_{v}L^{1}_{x}})$
    for some $p>\frac{3}{3-\gamma}$.
@@
$\mathcal{H}(\mu_{t})-\mathcal{H}(\mu_{s})=\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{12}}\tilde{\nabla}\log f_{r}{\,\rm d}\mathcal{U}_{r}{\,\rm d}r$ (3.9)
@@
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{12}}\tilde{\nabla}\log f_{t}{\,\rm d}\mathcal{U}_{t}$ (3.10)
@@
  • On the domain $\mathbb{T}^{3}\times\mathbb{R}^{3}$, we do not need to assume the moments bounds in $x$, since $\mathbb{T}^{3}$ can be treated as a bounded domain.
@@
$\partial_{t}\mu_{t}+\frac{1}{2}\tilde{\nabla}\cdot\mathcal{U}_{t}=0,\quad\mu\in\mathcal{P}(\mathbb{R}^{3}),\quad\mathcal{U}\in\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$
has been established in [carrillo2024landau]. In the proof of Theorem 3.6, we follow [carrillo2024landau] to treat the collision terms, where we do appropriate modification concerning the spatial variable $x\in\mathbb{R}^{3}$. We follow [EH24] to treat the transport part $v\cdot\nabla_{x}\mu_{t}$ that appears in the transport grazing rate equations (3.1). The second author and Erbar [EH24] consider a chain rule for a transport collision rate equation related to a delocalised fuzzy Boltzmann equation.
@@ 3.2.1 Regularisation
    Let $M(x)=c\exp\big{(}-\langle x\rangle\big{)}$, where we choose the constant $c>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{3})}=1$.
@@
    $M_{\alpha}(x)=c\alpha^{-3}\exp\big{(}-\langle\frac{x}{\alpha}\rangle\big{)}.$
@@
    For test functions $\phi:\mathbb{R}^{3}\times\mathbb{R}^{3}\to\mathbb{R}$ and $\Phi:\mathbb{R}^{12}\to\mathbb{R}$, we have
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\phi{\,\rm d}\mu^{\alpha}_{t}(x,v)=\int_{\mathbb{R}^{3\times 3}}\phi(x-y,v)M_{\alpha}(y){\,\rm d}\mu_{t}(x,v){\,\rm d}y,$
    $\int_{\mathbb{R}^{12}}\Phi{\,\rm d}\mathcal{U}^{\alpha}_{t}(x,x_{*},v,v_{*})$
    $=\int_{\mathbb{R}^{12}\times\mathbb{R}^{3}}\Phi(x-y,x_{*}-y_{*},v,v_{*})M_{\alpha}(y)M_{\alpha}(y_{*}){\,\rm d}\mathcal{U}_{t}(x,x_{*},v,v_{*}){\,\rm d}y{\,\rm d}y_{*}.$
@@
    where we choose $a>8+|\gamma|$ to ensure the integral bounds in Lemma 3.9 hold also for $g$, and we choose $Z>0$ such that $\|g\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}=1$.
@@
    We denote $f^{\alpha,\beta,\delta,\theta}_{t}$ and $U^{\alpha,\beta,\delta,\theta}_{t}$ the density of $\mu^{\alpha,\beta,\delta,\theta}_{t}$ and $\mathcal{U}^{\alpha,\beta,\delta,\theta}_{t}$ with respect to Lebesgue measure $\mathcal{L}$ on $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $\mathbb{R}^{12}$.
@@
  • The lower bound regularisation (3.11) is only needed in Case 2.4 in Assumption 2.1, where we assume the uniform moments bounds in $v$ and only time-integrable moments bounds in $x$
    $\sup_{t\in[0,T]}\mathcal{E}_{0,2}(\mu_{t})<+\infty\quad\text{and}\quad\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty.$
    The regularisation (3.12) provides a uniform-in-time positive lower bound for the regularised curve. Such bounds are needed for example in Lemma 3.9-3.
@@
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{t}{\,\rm d}x{\,\rm d}v=1$
@@
    for all $(t,x,v)\in I^{\delta}\times\mathbb{R}^{3}\times\mathbb{R}^{3}$, where the constant $C=C(\alpha,\beta)>0$.
@@
    $\int_{I^{\delta}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha,\beta,\delta}_{t}\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $=\int_{I^{\delta}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\delta}_{t}(y,u)\big{(}\int_{\mathbb{R}^{3}}M_{\alpha}(x-y){\,\rm d}x\big{)}\big{(}\int_{\mathbb{R}^{3}}M_{\beta}(v-u)\langle v\rangle^{a}{\,\rm d}v\big{)}{\,\rm d}y{\,\rm d}u{\,\rm d}t$
    $\leq C(\beta,a)\int_{0}^{T}\int_{-\delta}^{\delta}\eta^{\delta}_{s}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{t-s}(y,u)\langle u\rangle^{a}{\,\rm d}y{\,\rm d}u{\,\rm d}s{\,\rm d}t$
    $\leq C\int_{0}^{T}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f_{t}(x,v)\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
@@
    We show (3.15). We denote $F^{\beta,\delta}_{t}(v)=\int_{\mathbb{R}^{3}}f^{\beta,\delta}_{t}(x,v){\,\rm d}x$, then we have
    $\|f^{\alpha,\beta,\delta}_{t}\|_{L^{q}_{v}L^{1}_{x}}^{q}$ $=\int_{\mathbb{R}^{3}}\big{(}F^{\beta,\delta}(v)\big{)}^{q}{\,\rm d}v$
    $=\int_{\mathbb{R}^{3}}\Big{(}\int_{-\delta}^{\delta}\int_{\mathbb{R}^{3}}F_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s\Big{)}^{q}{\,\rm d}v$
    $\leq\int_{\mathbb{R}^{3}}\int_{-\delta}^{\delta}\int_{\mathbb{R}^{3}}F^{q}_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s{\,\rm d}v$
    $\leq\sup_{t\in[0,T]}\int_{\mathbb{R}^{3}}F^{q}_{t}(v){\,\rm d}v=\sup_{t\in[0,T]}\|f_{t}\|_{L^{q}_{v}L^{1}_{x}},$
@@
    $f^{\alpha,\beta,\delta}_{t}(x,v)$ $\leq\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.20)
    $\leq(\alpha\beta)^{-1}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\delta}{\,\rm d}x{\,\rm d}v=(\alpha\beta)^{-1}.$
@@
    We follow [carrillo2024landau, Lemma 30] with an appropriate modification for the spatial variable $x\in\mathbb{R}^{3}$. We note that
@@
    $f^{\alpha,\beta,\delta}(x,v)=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.22)
@@
    ×$\quad\times\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\delta}(y,w)\exp\Big{(}-\sqrt{2}\big{(}\langle\frac{w}{\beta}\rangle+\langle\frac{y}{\alpha}\rangle\big{)}\Big{)}{\,\rm d}y{\,\rm d}w$
@@
[…]
@@
    since $f_{t}\in\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ for all $t\in[0,T]$. Then there exist a time-independent $R_{1}$ and time-dependent $R_{2}(t)$ large, such that
@@
    $\Big{|}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha,\beta,\delta}_{t}\log f^{\alpha,\beta,\delta}_{t}\Big{|}$ $\leq C\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f^{\alpha,\beta,\delta}_{t}\big{(}\langle x\rangle+\langle v\rangle\big{)}{\,\rm d}x{\,\rm d}v$
@@
    where we use the moments bounds (3.13).
@@
$\mathcal{D}(f)=\frac{1}{2}\int_{\mathbb{R}^{12}}A\kappa ff_{*}\big{|}\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big{)}\big{|}^{2}{\,\rm d}\eta.$ (3.23)
We define ¯$\bar{\nabla}\mathop{=}\limits^{\textrm{def}}\nabla_{v}-\nabla_{v_{*}}$. The following identity holds
@@
[…]
@@
$\mathcal{D}(f)=2\int_{\mathbb{R}^{12}}A\kappa\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(ff_{*})\big{|}^{2}}{ff_{*}}{\,\rm d}\eta.$ (3.24)
We have the following estimates on regularised entropy dissipation and curve action. Notice that we only have the bounds on $\mathcal{D}(f^{\alpha,\delta})$ (with vanishing $\beta=0$), besides in the case in Section 4.
@@
$A\kappa\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(f^{\alpha}f^{\alpha}_{*})\big{|}^{2}}{f^{\alpha}f^{\alpha}_{*}}\leq A\kappa\Big{(}\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(ff_{*})\big{|}^{2}}{ff_{*}}\Big{)}*_{x,x_{*}}M_{\alpha}$ (3.25)
@@
$\mathcal{D}(f^{\alpha})=2\int_{\mathbb{R}^{12}}A\kappa\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(f^{\alpha}f^{\alpha}_{*})\big{|}^{2}}{f^{\alpha}f^{\alpha}_{*}}{\,\rm d}\eta.$
Notice that the regularisations and $\Pi_{(v-v_{*})^{\perp}}$ and $\bar{\nabla}$ are commuted. By using of the convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$ and Jensen’s inequality we have the pointwise bounds (3.25). By definition of the spatial kernel $\kappa$ in Assumption 2.1 (we refer to property (2.9) for Case 2.2 and definition (2.4) for Case 2.4), we have
@@
$\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t$ $\leq\int_{0}^{T}\int_{\mathbb{R}^{12}}A\big{(}M_{\alpha}*_{x,x_{*}}\kappa\big{)}\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(ff_{*})\big{|}^{2}}{ff_{*}}{\,\rm d}\eta{\,\rm d}t$
@@
[…]
@@
For any fixed $\alpha,\,\beta,\,\delta,\,\theta\geq 0$ and $\gamma\in(-3,1]$, we have
$\int_{I^{\delta}}\int_{\mathbb{R}^{12}}|U_{t}^{\alpha,\beta,\delta,\theta}|\big{(}\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big{)}{\,\rm d}\eta{\,\rm d}t<+\infty,$ (3.28)
@@
$\int_{I^{\delta}}\int_{\mathbb{R}^{12}}|U_{t}^{\alpha,\beta,\delta,\theta}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t<+\infty.$ (3.29)

### v2

@@ 3.2 Chain rule
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0.$
@@
Let $\gamma\in(-\gamma_{d},1]$. Let $(\mu_{t},\mathcal{U}_{t})\in\mathcal{TGRE}_{T}$ such that $|\mathcal{H}(\mu_{0})|<+\infty$ and $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{2d})$. We assume
@@
    for some $p>\frac{d}{d-|\gamma|}$.
@@
    $\text{and}\quad\langle v\rangle^{\gamma}f\in L^{\infty}([0,T];{L^{p}_{v}L^{1}_{x}})$
    for some $p>\frac{d}{d-\gamma}$.
@@
$\mathcal{H}(\mu_{t})-\mathcal{H}(\mu_{s})=\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log f_{r}{\,\rm d}\mathcal{U}_{r}{\,\rm d}r$ (3.9)
@@
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log f_{t}{\,\rm d}\mathcal{U}_{t}$ (3.10)
@@
  • On the domain $\mathbb{T}^{d}\times\mathbb{R}^{d}$, we do not need to assume the moment bounds in $x$, since $\mathbb{T}^{d}$ can be treated as a bounded domain.
@@
Remark 3.8 (Assumptions on curves).
  • The curve assumptions in Assumption 2.7 imply the curve assumptions in Theorem 3.6. More precisely, in the hard potential case $\gamma\in(0,1]$, in Assumption 2.7, we only assume moment and $L^{p}$-bounds, which in turn ensure the weighted $L^{p}$-bounds required in Theorem 3.6. Let
    $f\in L^{\infty}([0,T];L^{1}_{0,2+\gamma}(\mathbb{R}^{2d}))\cap L^{\infty}([0,T];L^{p}_{v}L^{1}_{x})$
    for some $\frac{d}{d-\gamma}<p$, then by Hölder inequality, we have
    $\|\langle v\rangle^{|\gamma|}f\|_{L^{p^{\prime}}_{v}L^{1}_{x}}\leq\|f\|_{L^{p}_{v}L^{1}_{x}}^{\frac{p-p^{\prime}}{p}}\|f\|_{L^{1}_{0,2+|\gamma|}(\mathbb{R}^{2d})}^{\frac{p^{\prime}}{p}},$
    and hence,
    $\langle v\rangle^{\gamma}f\in L^{\infty}([0,T];L^{p^{\prime}}_{v}L^{1}_{x})$
    for all $\max\big(\frac{d}{d-\gamma},\frac{p\gamma}{2+\gamma}\big)<p^{\prime}<p$.
  • There are $L^{p}_{v}L^{1}_{x}$ assumptions on curves in both the soft and hard potential cases, though they play different roles. In the hard potential case $\gamma\in(0,1]$, the $L^{p}$-assumptions are used to bound the weighted Fisher information, see Lemma 2.9. In the soft potential case $\gamma\in(-\gamma_{d},0)$, the $L^{p}$-assumptions are used to control $\|\langle v\rangle^{2}f*_{v}|v|^{\gamma}\|_{L^{\infty}_{v}L^{1}_{x}}$, see, for example, (3.45).
  • Theorem 3.6 holds, if we replace $L^{p}_{v}L^{1}_{x}$ in the assumptions by $L^{1}_{x}L^{p}_{v}$. The $L^{p}$ bound is used to ensure the integrability of $\int_{\mathbb{R}^{4d}}ff_{*}|v-v_{*}|^{\gamma}{\,\rm d}\eta$, for which we use Lemma 2.10, and we make no distinction between $L^{1}_{x}L^{p}_{v}$ and $L^{p}_{v}L^{1}_{x}$ there.
@@
$\partial_{t}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0,\quad\mu\in\mathcal{P}(\mathbb{R}^{d}),\quad\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$
has been established in [CDDW24]. In the proof of Theorem 3.6, we follow [CDDW24] to treat the collision terms, where we do appropriate modification concerning the spatial variable $x\in\mathbb{R}^{d}$. We follow [EH25] to treat the transport part $v\cdot\nabla_{x}\mu_{t}$ that appears in the transport grazing rate equations (3.1). The second author and Erbar [EH25] consider a chain rule for a transport collision rate equation related to a delocalised fuzzy Boltzmann equation.
@@ 3.2.1 Regularisation
    Let $M(x)=c\exp\big(-\langle x\rangle\big)$, where we choose the constant $c>0$ such that $\|M\|_{L^{1}(\mathbb{R}^{d})}=1$.
@@
    $M_{\alpha}(x)=c\alpha^{-d}\exp\big(-\langle\frac{x}{\alpha}\rangle\big).$
@@
    For test functions $\phi:\mathbb{R}^{2d}\to\mathbb{R}$ and $\Phi:\mathbb{R}^{4d}\to\mathbb{R}$, we have
    $\int_{\mathbb{R}^{2d}}\phi{\,\rm d}\mu^{\alpha}_{t}(x,v)=\int_{\mathbb{R}^{3d}}\phi(x-y,v)M_{\alpha}(y){\,\rm d}\mu_{t}(x,v){\,\rm d}y,$
    $\int_{\mathbb{R}^{4d}}\Phi{\,\rm d}\mathcal{U}^{\alpha}_{t}(x,x_{*},v,v_{*})$
    $=\int_{\mathbb{R}^{6d}}\Phi(x-y,x_{*}-y_{*},v,v_{*})M_{\alpha}(y)M_{\alpha}(y_{*}){\,\rm d}\mathcal{U}_{t}(x,x_{*},v,v_{*}){\,\rm d}y{\,\rm d}y_{*}.$
@@
    where we choose $a>8+|\gamma|$ to ensure the integral bounds in Lemma 3.10 hold also for $g$, and we choose $Z>0$ such that $\|g\|_{L^{1}(\mathbb{R}^{2d})}=1$.
@@
    We denote $f^{\alpha,\beta,\delta,\theta}_{t}$ and $U^{\alpha,\beta,\delta,\theta}_{t}$ the density of $\mu^{\alpha,\beta,\delta,\theta}_{t}$ and $\mathcal{U}^{\alpha,\beta,\delta,\theta}_{t}$ with respect to Lebesgue measure $\mathcal{L}$ on $\mathbb{R}^{2d}$ and $\mathbb{R}^{4d}$.
@@
  • We use the lower bound regularisation (3.11) to provide a uniform-in-time positive lower bound for the regularised curve, in particular in $x$, since we only time-integrable moment bounds in $x$
    $\int_{0}^{T}\mathcal{E}_{2,0}(\mu_{t}){\,\rm d}t<+\infty.$
    The uniform lower bounds are needed, for example, in Lemma 3.10-3.
@@
    $\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta,\theta}_{t}{\,\rm d}x{\,\rm d}v=\int_{\mathbb{R}^{2d}}f_{t}{\,\rm d}x{\,\rm d}v=1$
@@
    for all $(t,x,v)\in I^{\delta}\times\mathbb{R}^{2d}$, where the constant $C=C(\alpha,\beta)>0$.
@@
    $\int_{I^{\delta}}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t$
    $=\int_{I^{\delta}}\int_{\mathbb{R}^{2d}}f^{\delta}_{t}(y,u)\big(\int_{\mathbb{R}^{d}}M_{\alpha}(x-y){\,\rm d}x\big)\big(\int_{\mathbb{R}^{d}}M_{\beta}(v-u)\langle v\rangle^{a}{\,\rm d}v\big){\,\rm d}y{\,\rm d}u{\,\rm d}t$
    $\leq C(\beta,a)\int_{0}^{T}\int_{-\delta}^{\delta}\eta^{\delta}_{s}\int_{\mathbb{R}^{2d}}f_{t-s}(y,u)\langle u\rangle^{a}{\,\rm d}y{\,\rm d}u{\,\rm d}s{\,\rm d}t$
    $\leq C\int_{0}^{T}\int_{\mathbb{R}^{2d}}f_{t}(x,v)\langle v\rangle^{a}{\,\rm d}x{\,\rm d}v{\,\rm d}t,$
@@
    We show (3.15). We denote $F^{\beta,\delta}_{t}(v)=\int_{\mathbb{R}^{d}}f^{\beta,\delta}_{t}(x,v){\,\rm d}x$, then we have
    $\|f^{\alpha,\beta,\delta}_{t}\|_{L^{q}_{v}L^{1}_{x}}^{q}$ $=\int_{\mathbb{R}^{d}}\big(F^{\beta,\delta}(v)\big)^{q}{\,\rm d}v$
    $=\int_{\mathbb{R}^{d}}\Big(\int_{-\delta}^{\delta}\int_{\mathbb{R}^{d}}F_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s\Big)^{q}{\,\rm d}v$
    $\leq\int_{\mathbb{R}^{d}}\int_{-\delta}^{\delta}\int_{\mathbb{R}^{d}}F^{q}_{t-s}(u)M_{\beta}(v-u)\eta^{\delta}_{s}{\,\rm d}u{\,\rm d}s{\,\rm d}v$
    $\leq\sup_{t\in[0,T]}\int_{\mathbb{R}^{d}}F^{q}_{t}(v){\,\rm d}v=\sup_{t\in[0,T]}\|f_{t}\|_{L^{q}_{v}L^{1}_{x}},$
@@
    $f^{\alpha,\beta,\delta}_{t}(x,v)$ $\leq\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.20)
    $\leq(\alpha\beta)^{-1}\int_{\mathbb{R}^{2d}}f^{\delta}{\,\rm d}x{\,\rm d}v=(\alpha\beta)^{-1}.$
@@
    We follow [CDDW24, Lemma 30] with an appropriate modification for the spatial variable $x\in\mathbb{R}^{d}$. We note that
@@
    $f^{\alpha,\beta,\delta}(x,v)=\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)M_{\beta}(v-w)M_{\alpha}(x-y){\,\rm d}y{\,\rm d}w$ (3.22)
@@
    $\quad\times\int_{\mathbb{R}^{2d}}f^{\delta}(y,w)\exp\Big(-\sqrt{2}\big(\langle\frac{w}{\beta}\rangle+\langle\frac{y}{\alpha}\rangle\big)\Big){\,\rm d}y{\,\rm d}w$
@@
    $\leq{}$ $\int_{\{|v|\geq R_{1}\}\times\mathbb{R}^{d}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v+\int_{\{|x|\geq R_{2}\}\times\mathbb{R}^{d}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v$
    $\leq{}$ $\int_{\{|v|\geq R_{1}\}\times\mathbb{R}^{d}}\frac{|v|^{2}}{R_{1}^{2}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v+\int_{\{|x|\geq R_{2}\}\times\mathbb{R}^{d}}\frac{|x|^{2}}{R_{2}^{2}}f^{\delta}(x,v){\,\rm d}x{\,\rm d}v$
    $\leq{}$ $\frac{\mathcal{E}_{0,2}(f^{\delta}_{t})}{R_{1}^{2}}+\frac{\mathcal{E}_{2,0}(f^{\delta}_{t})}{R_{2}^{2}}.$
@@
    since $f_{t}\in\mathcal{P}_{2,2+|\gamma|}(\mathbb{R}^{2d})$ for all $t\in[0,T]$. Then there exist a time-independent $R_{1}$ and time-dependent $R_{2}(t)$ large, such that
@@
    $\Big|\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\log f^{\alpha,\beta,\delta}_{t}\Big|$ $\leq C\mathcal{E}_{2,0}(f^{\delta}_{t})^{\frac{1}{2}}\int_{\mathbb{R}^{2d}}f^{\alpha,\beta,\delta}_{t}\big(\langle x\rangle+\langle v\rangle\big){\,\rm d}x{\,\rm d}v$
@@
    where we use the moment bounds (3.13).
@@
$\mathcal{D}(f)=\frac{1}{2}\int_{\mathbb{R}^{4d}}A\kappa ff_{*}\big|\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}\log f-\nabla_{v_{*}}\log f_{*}\big)\big|^{2}{\,\rm d}\eta.$ (3.23)
We define $\overline{\nabla}\mathop{=}\limits^{\textrm{def}}\nabla_{v}-\nabla_{v_{*}}$. The following identity holds
@@
$={}$ $4\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}\sqrt{ff_{*}}\big|^{2}=4\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}.$
@@
$\mathcal{D}(f)=2\int_{\mathbb{R}^{4d}}A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}{\,\rm d}\eta.$ (3.24)
We have the following estimates on regularised entropy dissipation and curve action. Notice that we only have the bounds on $\mathcal{D}(f^{\alpha,\delta})$ (with vanishing $\beta=0$), besides in the case in Section 4.1.
@@
$A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(f^{\alpha}f^{\alpha}_{*})\big|^{2}}{f^{\alpha}f^{\alpha}_{*}}\leq A\kappa\Big(\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}\Big)*_{x,x_{*}}M_{\alpha}$ (3.25)
@@
$\mathcal{D}(f^{\alpha})=2\int_{\mathbb{R}^{4d}}A\kappa\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(f^{\alpha}f^{\alpha}_{*})\big|^{2}}{f^{\alpha}f^{\alpha}_{*}}{\,\rm d}\eta.$
Notice that the regularisations and $\Pi_{(v-v_{*})^{\perp}}$ and $\overline{\nabla}$ are commuted. By using of the convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$ and Jensen’s inequality we have the pointwise bounds (3.25). By definition of the spatial kernel $\kappa$ in Assumption 2.1 and in Remark 2.2 (we refer to property (2.10) and definition (2.4)), we have
@@
$\int_{0}^{T}\mathcal{D}(f^{\alpha}){\,\rm d}t$ $\leq\int_{0}^{T}\int_{\mathbb{R}^{4d}}A\big(M_{\alpha}*_{x,x_{*}}\kappa\big)\frac{\big|\Pi_{(v-v_{*})^{\perp}}\overline{\nabla}(ff_{*})\big|^{2}}{ff_{*}}{\,\rm d}\eta{\,\rm d}t$
@@
$\leq{}$ $\int_{0}^{T}\int_{\mathbb{R}^{4d}}M_{\beta}*_{v,v_{*}}\big(M_{\alpha}*_{x,x_{*}}\kappa^{-1}\big)\frac{|U_{t}|^{2}}{ff_{*}}{\,\rm d}\eta{\,\rm d}t$
$\leq{}$ $C\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{\kappa ff_{*}}{\,\rm d}\eta{\,\rm d}t=C\int_{0}^{T}\mathcal{A}(f,U){\,\rm d}t.$
@@
For any fixed $\alpha,\,\beta,\,\delta,\,\theta\geq 0$ and $\gamma\in(-\gamma_{d},1]$, we have
$\int_{I^{\delta}}\int_{\mathbb{R}^{4d}}|U_{t}^{\alpha,\beta,\delta,\theta}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big){\,\rm d}\eta{\,\rm d}t<+\infty,$ (3.28)
@@
$\int_{I^{\delta}}\int_{\mathbb{R}^{4d}}|U_{t}^{\alpha,\beta,\delta,\theta}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t<+\infty.$ (3.29)